Chapter Section , ustion 17 Stics becobonded to look for
Answer #1

• Solution : As i can see in the picture 0 - we have given differential equation y (2d-15)yt (2² 152+50)y=0 - We need to fin= (22-15) 15 2 so the roots of equation mi = 22 +10 = 2-5 ® - - S m₂ = 22-20=2-10 ® are so, the solutions of equation (2-5)+9.2 PPP similarly for Y o as to ipfo 2-1020 le 2210 =) de 100, 10) - € Therefore from equation ② auf 6 26(-0,5) Hem if take e